The sum of three consecutive odd integers is 201. Find the integers.
Answer by stanbon(75887)

The sum of three consecutive odd integers is 201. Find the integers.
------------
1st odd: 2x+1
2nd odd; 2x+3
3rd odd: 2x+5
-----------------
Add to get: 6x + 9 = 201
6x = 192
x = 32
---------
2x+1 = 65
2x+3 = 67
2x+5 = 69
=============
Cheers,
Stan H.